# Gradient-based Sample Selection for Faster Bayesian Optimization

### v2

@@ 6.2 Real-World Application
To assess the applicability of GSSBO in real-world applications, we applied NAS to a diabetes-detection problem from the UCI repository [47]. We modeled the problem of searching for optimal hyperparameters as a BO problem. Specifically, each query $(x_{t},y_{t})$ corresponds to a choice of ($[32,128],[1e-6,1.0],[1e-6,1.0],[1,8]$ for batch size, Learning Rate, Learning Rate decay, hidden dim), where $y_{t}$ is the test classification error. Each iteration is repeated 5 times. The results in Figure 2 indicate that GSSBO outperforms all competitors.
@@ 7 Conclusion
BO is known to be effective for optimization in settings where the objective function is expensive to evaluate. In large-budget scenarios, the use of a full GP model can slow the convergence of BO, leading to poor scaling in these cases. In this paper, we investigated the use of gradient-based sample selection to accelerate BO, demonstrating how a carefully constructed subset, guided by gradient information, can serve as an efficient surrogate for the full sample set, significantly enhancing the efficiency of the BO process. As we have shown in a comprehensive set of experiments, the proposed GSSBO shows its ability to significantly reduce computational time while maintaining competitive optimization performance. Synthetic benchmarks highlight its scalability across various problem dimensions, while real-world applications confirm its practical utility. The sensitivity analysis further showcases the adaptability of the method to different parameter settings, reinforcing its robustness in diverse optimization scenarios. Overall, these findings underline the potential of gradient-based sample selection in addressing the scaling challenges of BO.
@@ A.1 Theorem 1:Analysis on subset GP:
Define the following positive definite matrices: $\mathbf{M_{A}}=\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I},~{}\mathbf{M_{B}}=\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}.$ Then, the difference between $\mathbf{M_{A}}$ and $\mathbf{M_{B}}$ is: $\mathbf{M_{A}}-\mathbf{M_{B}}=\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}.$ We have $\|\mathbf{M_{A}}^{-1}-\mathbf{M_{B}}^{-1}\|\;\leq\;\|\mathbf{M_{A}}^{-1}\|\;\|\mathbf{M_{B}}^{-1}\|\;\|\mathbf{M_{A}}-\mathbf{M_{B}}\|,$ and take $C_{M}\;=\;\|\mathbf{M_{A}}^{-1}\|\;\|\mathbf{M_{B}}^{-1}\|.$ Since $\mathbf{M_{A}},\mathbf{M_{B}}$ are positive definite, denote $\lambda_{\min}(\mathbf{M_{A}})=\text{the smallest eigenvalue of }\mathbf{M_{A}},~{}\lambda_{\min}(\mathbf{M_{B}})=\text{the smallest eigenvalue of }\mathbf{M_{B}}.$ Under the spectral norm, $\|\mathbf{M_{A}}^{-1}\|=\frac{1}{\lambda_{\min}(\mathbf{M_{A}})},\|\mathbf{M_{B}}^{-1}\|=\frac{1}{\lambda_{\min}(\mathbf{M_{B}})}.$ Noting that $\lambda_{\min}(\mathbf{M_{A}})=\lambda_{\min}(K_{\mathcal{D}\mathcal{D}})+\sigma_{n}^{2},~{}\lambda_{\min}(\mathbf{M_{B}})\geq\sigma_{n}^{2},$ we obtain the bound on $C_{M}$ that $C_{M}=\frac{1}{\bigl{(}\lambda_{\min}(K_{\mathcal{D}\mathcal{D}})+\sigma_{n}^{2}\bigr{)}\,\sigma_{n}^{2}}\leq\frac{1}{\lambda_{\min}(K_{\mathcal{D}\mathcal{D}})\,\sigma_{n}^{2}}.$ We have: $\|(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|\leq C_{M}\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.$ Substituting this into the expression for $\Delta\mu(\mathbf{x}_{*})$, we obtain:
$\begin{split}|\Delta\mu(\mathbf{x}_{*})|&\leq\|\mathbf{k}_{*\mathcal{D}}\|\;\|\mathbf{y}\|\;\bigl{\|}(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\bigr{\|}\\[4.30554pt] &\leq\|\mathbf{k}_{*\mathcal{D}}\|\;\|\mathbf{y}\|\;C_{M}\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.\end{split}$ (7)
@@
$\begin{split}\bigl{|}\Delta\sigma^{2}(\mathbf{x}_{*})\bigr{|}&\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}\,\Bigl{\|}(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\Bigr{\|}\\ &\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}\,C_{M}\,\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.\end{split}$ (9)
@@
Now we force the initial batch of $n_{0}$ samples $\{\mathbf{x}_{1},\dots,\mathbf{x}_{n_{0}}\}$ and the latest sample $\mathbf{x}_{\mathrm{L}}$ to be included in $\mathcal{U}$. We then choose the remaining $M-n_{0}-1$ samples to maximize gradient variance. We select the $M-n_{0}-1$ samples that maximize the variance of gradient information. We have $\mathcal{U}\;=\;\{\mathbf{x}_{1},\dots,\mathbf{x}_{n_{0}}\}\;\cup\;\{\mathbf{x}_{\mathrm{L}}\}\;\cup\;\mathcal{U}^{\prime},$ where $|\mathcal{U}^{\prime}|=M-n_{0}-1$ so that $|\mathcal{U}|=M$, the total subset $\mathcal{U}$ is still of size $M$. We consider a subset $\mathcal{U}\subset\mathcal{D}$ of size $M<N$ to build a low-rank approximation of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ are derived from $\mathcal{U}$. Let $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ be a positive definite kernel matrix with eigen-decomposition: $\mathbf{K}_{\mathcal{D}\mathcal{D}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ where $\bm{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{N}\geq 0$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{N}]$. The best rank-$M$ approximation to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ in spectral norm is: $\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}=\mathbf{U}_{M}\bm{\Lambda}_{M}\mathbf{U}_{M}^{\top},$ where $\mathbf{U}_{M}=[\mathbf{u}_{1},\dots,\mathbf{u}_{M}]$ and $\bm{\Lambda}_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. By the Eckart–Young–Mirsky theorem: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|=\lambda_{M+1}.$ Suppose $\mathcal{U}\subset\mathcal{D}$, $|\mathcal{U}|=M$, produces a approximation: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$
Gradient-based select $\mathcal{U}$. Here, we give a derivation showing that selecting the subset $\mathcal{U}$ by maximizing gradient-direction diversity is equivalent to the standard greedy Nyström column-selection. For a kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ and standard basis vectors $\bm{e}_{i}\in\mathbb{R}^{N}$, define $\bm{\phi}_{i}=\mathbf{K}_{\mathcal{D}\mathcal{D}}\,\bm{e}_{i}\quad\text{(``kernel column" for sample $i$)},\qquad\bm{\psi}_{i}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\bm{e}_{i}\quad\text{(``gradient column" for sample $i$)}.$ Suppose after $t$ steps we have selected index set $\mathcal{U}_{t}\subset\{1,\dots,N\}$. Let $\mathbf{C}_{t}=[\,\bm{\phi}_{j}\,]_{j\in\mathcal{U}_{t}}\in\mathbb{R}^{N\times t},\qquad\mathbf{G}_{t}=[\,\bm{\psi}_{j}\,]_{j\in\mathcal{U}_{t}}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\mathbf{C}_{t}.$ Define the orthogonal projectors $\mathbf{P}_{t}^{(K)}=\mathbf{C}_{t}\,(\mathbf{C}_{t}^{\top}\mathbf{C}_{t})^{-1}\,\mathbf{C}_{t}^{\top},\qquad\mathbf{P}_{t}^{(g)}=\mathbf{G}_{t}\,(\mathbf{G}_{t}^{\top}\mathbf{G}_{t})^{-1}\,\mathbf{G}_{t}^{\top}.$ By construction, $\mathrm{range}(\mathbf{P}_{t}^{(K)})=\mathrm{range}(\mathbf{P}_{t}^{(g)})=\mathrm{span}\{\mathbf{C}_{t}\}.$ Using $\mathbf{G}_{t}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\mathbf{C}_{t}$ one checks $\mathbf{P}_{t}^{(g)}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\mathbf{C}_{t}\,\bigl{(}\mathbf{C}_{t}^{\top}\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-2}\mathbf{C}_{t}\bigr{)}^{-1}\,\mathbf{C}_{t}^{\top}\,\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1},$ and one can verify the key identity $\mathbf{I}-\mathbf{P}_{t}^{(g)}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\,\bigl{(}\mathbf{I}-\mathbf{P}_{t}^{(K)}\bigr{)}\,\mathbf{K}_{\mathcal{D}\mathcal{D}}.$ At iteration $t+1$: Gradient-diversity rule picks $i_{t+1}=\arg\max_{i\notin\mathcal{U}_{t}}\bigl{\lVert}(\mathbf{I}-\mathbf{P}_{t}^{(g)})\,\bm{\psi}_{i}\bigr{\rVert}.$ Nyström column-selection rule picks $i_{t+1}=\arg\max_{i\notin\mathcal{U}_{t}}\bigl{\lVert}(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\bigr{\rVert}.$ Using $\bm{\psi}_{i}=\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}\bm{e}_{i}$ and the conjugation relation, $\bigl{\lVert}(\mathbf{I}-\mathbf{P}_{t}^{(g)})\,\bm{\psi}_{i}\bigr{\rVert}=\bigl{\lVert}\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\bigr{\rVert}.$ Since $\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}$ is a fixed invertible operator, maximizing $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\|$ is equivalent to maximizing $\|(\mathbf{I}-\mathbf{P}_{t}^{(K)})\,\bm{\phi}_{i}\|$. Therefore both rules select the same index $i_{t+1}$ at every step. Thus, selecting $\mathcal{U}$ by maximizing gradient-direction diversity coincides with the greedy Nyström column-selection.
@@
If we choose $\mathcal{U}$ via a greedy procedure that maximizes gradient-information, then we defines the orthogonal projector onto the selected columns by $\mathbf{P_{\mathcal{U}}}\;=\;\mathbf{K_{D\,U}\,(K_{U\,U})^{-1}K_{U\,D}},$ and measures the subspace approximation error against the true principal eigenspace $\mathrm{span}(U_{M})$ as
@@
Random select $\mathcal{U}$. If the subset $\mathcal{U}$ is chosen to approximate the principal eigenspace spanned by randomly choose $\mathbf{U}_{M}$, then Nyström approximation theory [48] guarantees that: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\leq\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}=\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}.$ where $\epsilon$ is an error control parameter related to the number of columns sampled randomly in the approximation. In particular, if one samples $c$ columns uniformly at random , then for any $\delta\in(0,1)$ the following holds with probability at least $1-\delta$ [48]:
@@
Comparison. Let $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ have eigenpairs $(\lambda_{k},u_{k})$, and let $U_{M}=[\,u_{1},\dots,u_{M}\,]\in\mathbb{R}^{N\times M}$ be the matrix of its top $M$ eigenvectors. Suppose we build a Nyström approximation by selecting columns greedily in $M$ steps; let $\mathcal{U}_{t}$ be the chosen indices after $t$ steps and $\mathbf{P_{t}}$ the corresponding orthogonal projector onto the span of those columns. Define the residual energy $E_{t}\;=\;\sum_{k=1}^{M}\|(\mathbf{I}-\mathbf{P_{t}})\mathbf{u_{k}}\|^{2}.$ Observe $\sum_{i=1}^{N}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}=\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{K}\|_{F}^{2}\;\geq\;\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\Lambda_{M}^{1/2}\|_{F}^{2}\;\geq\;\lambda_{1}\,\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\|_{F}^{2}=E_{t},$ where $\mathbf{\phi_{i}}=\mathbf{K}\,\mathbf{e_{i}}$ and $\Lambda_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. Hence among the remaining $N-t$ columns there is some index $j\notin\mathcal{U}_{t}$ with $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\geq\;\frac{1}{N-t}\sum_{i\notin\mathcal{U}_{t}}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}\;\geq\;\frac{E_{t}}{N-t}.$ Since the greedy rule picks precisely the column that maximizes $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|$, the new residual satisfies $E_{t+1}=E_{t}\;-\;\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\leq\;E_{t}\Bigl{(}1-\tfrac{1}{N-t}\Bigr{)}.$ Iterating the one-step bound from $t=0$ to $t=M-1$ gives $E_{M}\;\leq\;E_{0}\prod_{t=0}^{M-1}\Bigl{(}1-\tfrac{1}{N-t}\Bigr{)}=\;M\;\prod_{k=N-M+1}^{N}\Bigl{(}1-\tfrac{1}{k}\Bigr{)},$ since $E_{0}=\|U_{M}\|_{F}^{2}=M$. Using $1-\tfrac{1}{k}\leq e^{-1/k}$, we obtain $E_{M}\;\leq\;M\;\exp\!\Bigl{(}-\sum_{k=N-M+1}^{N}\tfrac{1}{k}\Bigr{)}\;\leq\;M\;\exp\!\bigl{(}-\tfrac{M}{N}\bigr{)}.$ Thus the final projection error $\epsilon_{g}=\sqrt{E_{M}}\;\leq\;\sqrt{M}\,\exp\!\bigl{(}-\tfrac{M}{2N}\bigr{)}=O\!\bigl{(}e^{-\tfrac{M}{2N}}\bigr{)}$. Because $\epsilon_{g}$ decays exponentially in $M$, it is asymptotically much smaller than the $O(1/\sqrt{M})$ decay of the random-sampling parameter $\epsilon$. Hence, greedy (gradient-based) selection yields a Nyström approximation error that is provably tighter than uniform random sampling for the same $M$.
The key insight is that each greedy step removes at least the average of the remaining projection error (or “energy”). If at step $t$ the total unexplained energy is $E_{t}$, then among the $N-t$ remaining columns there must be one whose residual contributes at least $E_{t}/(N-t)$. By selecting that column, the algorithm shrinks the leftover energy by a factor of $\bigl{(}1-\tfrac{1}{N-t}\bigr{)}$. Repeating this multiplicative reduction $M$ times drives the energy down by a product of $\bigl{(}1-\tfrac{1}{k}\bigr{)}$ terms, which is bounded by an exponential $e^{-M/N}$. Hence the projection error $\epsilon_{g}$ decays exponentially in $M$.
@@
$\begin{split}\bigl{|}\Delta\mu(\mathbf{x}_{*})\bigr{|}&\leq\|\mathbf{k}_{*\mathcal{D}}\|\,\|\mathbf{y}\|\,C_{M}\,\bigl{\|}\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\bigr{\|},\\ \bigl{|}\Delta\sigma(\mathbf{x}_{*})\bigr{|}&\leq\|\mathbf{k}_{*\mathcal{D}}\|\,\sqrt{C_{M}\,\bigl{\|}\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\bigr{\|}}.\end{split}$ (13)
@@
Here, $A_{n},B_{n}$ are determined by the number of iterations $n$, the same below. We now absorb both terms into a single, $n$–dependent penalty. Define $\sigma_{\min}(n)\;=\;\min_{x\in\mathcal{X}}\,{\sigma}_{n}(x),\qquad\delta_{n}\;=\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)},\qquad\tilde{\beta}_{n}\;=\;\beta_{n}+\delta_{n}.$ Since for all $x$, ${\sigma}_{n}(x)\geq\sigma_{\min}(n)>0$, it follows that $A_{n}+\beta_{n}\,B_{n}=\;\delta_{n}\,\sigma_{\min}(n)\;\leq\;\delta_{n}\,{\sigma}_{n}(x),$ and hence
@@
where $\tilde{\beta}_{n}=\;\beta_{n}+\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)}.$ This completes the single-penalty construction.

### v3

@@ 6.2 Real-World Application
To assess the applicability of GSSBO in real-world applications, we use a diabetes-detection problem from the UCI repository [Dua and Graff, 2017]. We modeled the problem of searching for optimal hyperparameters as a BO problem. Specifically, each query $(x_{t},y_{t})$ corresponds to a choice of ([32,128],[1e6,1.0],[1e6,1.0],[1,8][32,128],[1e-6,1.0],[1e-6,1.0],[1,8] for batch size, Learning Rate, Learning Rate decay, hidden dim), where $y_{t}$ is the test classification error. The results in Figure 2 indicate that GSSBO outperforms all competitors.
@@ 7 Conclusion
BO is known to be effective for optimization in settings where the objective function is expensive to evaluate. In large-budget scenarios, the use of a full GP model can slow the convergence of BO, leading to poor scaling in these cases. In this paper, we investigated the use of gradient-based sample selection to accelerate BO, demonstrating how a carefully constructed subset, guided by gradient information, can serve as an efficient surrogate for the full sample set, significantly enhancing the efficiency of the BO process. As we have shown in a comprehensive set of experiments, the proposed GSSBO shows its ability to significantly reduce computational time while maintaining competitive optimization performance. Synthetic and real-world benchmarks highlight its scalability and practical utility across various problem dimensions. The sensitivity analysis further showcases the adaptability of the method to different parameter settings. Overall, these findings underline the potential of gradient-based sample selection in addressing the BO scaling challenges.
Appendix
@@ A.1 Theorem 1:Analysis on subset GP:
[…]
@@
[…]
@@
Now we force the latest sample $\mathbf{x}_{\mathrm{L}}$ to be included in $\mathcal{U}$. We then choose the remaining M1M-1 samples to maximize gradient variance. We select the M1M-1 samples that maximize the variance of gradient information. We have $\mathcal{U}\;=\;\{\mathbf{x}_{\mathrm{L}}\}\;\cup\;\mathcal{U}^{\prime},$ where $|\mathcal{U}^{\prime}|=M-1$ so that $|\mathcal{U}|=M$, the total subset $\mathcal{U}$ is still of size $M$. We consider a subset $\mathcal{U}\subset\mathcal{D}$ of size $M<N$ to build a low-rank approximation of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ are derived from $\mathcal{U}$. Let $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ be a positive definite kernel matrix with eigen-decomposition: $\mathbf{K}_{\mathcal{D}\mathcal{D}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ where $\bm{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{N}\geq 0$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{N}]$. The best rank-$M$ approximation to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ in spectral norm is: $\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}=\mathbf{U}_{M}\bm{\Lambda}_{M}\mathbf{U}_{M}^{\top},$ where $\mathbf{U}_{M}=[\mathbf{u}_{1},\dots,\mathbf{u}_{M}]$ and $\bm{\Lambda}_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. By the Eckart–Young–Mirsky theorem: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|=\lambda_{M+1}.$ Suppose $\mathcal{U}\subset\mathcal{D}$, $|\mathcal{U}|=M$, produces a approximation: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$
[…]
@@
If we choose $\mathcal{U}$ via a greedy procedure that maximizes gradient‐information, then we defines the orthogonal projector onto the selected columns by $\mathbf{P_{\mathcal{U}}}\;=\;\mathbf{K_{D\,U}\,(K_{U\,U})^{-1}K_{U\,D}},$ and measures the subspace approximation error against the true principal eigenspace $\mathrm{span}(U_{M})$ as
@@
Random select $\mathcal{U}$. If the subset $\mathcal{U}$ is chosen to approximate the principal eigenspace spanned by randomly choose $\mathbf{U}_{M}$, then Nyström approximation theory [Drineas et al., 2005] guarantees that: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\leq\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}=\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}.$ where $\epsilon$ is an error control parameter related to the number of columns sampled randomly in the approximation. In particular, if one samples $c$ columns uniformly at random , then for any $\delta\in(0,1)$ the following holds with probability at least $1-\delta$ [Drineas et al., 2005]:
@@
Comparison. Let $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ have eigenpairs $(\lambda_{k},u_{k})$, and let $U_{M}=[\,u_{1},\dots,u_{M}\,]\in\mathbb{R}^{N\times M}$ be the matrix of its top $M$ eigenvectors. Suppose we build a Nyström approximation by selecting columns greedily in $M$ steps; let $\mathcal{U}_{t}$ be the chosen indices after $t$ steps and $\mathbf{P_{t}}$ the corresponding orthogonal projector onto the span of those columns. Define the residual energy $E_{t}\;=\;\sum_{k=1}^{M}\|(\mathbf{I}-\mathbf{P_{t}})\mathbf{u_{k}}\|^{2}.$ Observe $\sum_{i=1}^{N}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}=\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{K}\|_{F}^{2}\;\geq\;\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\Lambda_{M}^{1/2}\|_{F}^{2}\;\geq\;\lambda_{1}\,\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\|_{F}^{2}=E_{t},$ where $\mathbf{\phi_{i}}=\mathbf{K}\,\mathbf{e_{i}}$ and $\Lambda_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. Hence among the remaining NtN-t columns there is some index $j\notin\mathcal{U}_{t}$ with $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\geq\;\frac{1}{N-t}\sum_{i\notin\mathcal{U}_{t}}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}\;\geq\;\frac{E_{t}}{N-t}.$ Since the greedy rule picks precisely the column that maximizes $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|$, the new residual satisfies $E_{t+1}=E_{t}\;-\;\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\leq\;E_{t}\Bigl(1-\tfrac{1}{N-t}\Bigr).$ Iterating the one‐step bound from $t=0$ to t=M1t=M-1 gives $E_{M}\;\leq\;E_{0}\prod_{t=0}^{M-1}\Bigl(1-\tfrac{1}{N-t}\Bigr)=\;M\;\prod_{k=N-M+1}^{N}\Bigl(1-\tfrac{1}{k}\Bigr),$ since $E_{0}=\|U_{M}\|_{F}^{2}=M$. Using $1-\tfrac{1}{k}\leq e^{-1/k}$, we obtain $E_{M}\;\leq\;M\;\exp\!\Bigl(-\sum_{k=N-M+1}^{N}\tfrac{1}{k}\Bigr)\;\leq\;M\;\exp\!\bigl(-\tfrac{M}{N}\bigr).$ Thus the final projection error $\epsilon_{g}=\sqrt{E_{M}}\;\leq\;\sqrt{M}\,\exp\!\bigl(-\tfrac{M}{2N}\bigr)=O\!\bigl(e^{-\tfrac{M}{2N}}\bigr)$. Because $\epsilon_{g}$ decays exponentially in $M$, it is asymptotically much smaller than the $O(1/\sqrt{M})$ decay of the random‐sampling parameter $\epsilon$. Hence, greedy (gradient‐based) selection yields a Nyström approximation error that is provably tighter than uniform random sampling for the same $M$.
The key insight is that each greedy step removes at least the average of the remaining projection error (or “energy”). If at step $t$ the total unexplained energy is $E_{t}$, then among the NtN-t remaining columns there must be one whose residual contributes at least $E_{t}/(N-t)$. By selecting that column, the algorithm shrinks the leftover energy by a factor of $\bigl(1-\tfrac{1}{N-t}\bigr)$. Repeating this multiplicative reduction $M$ times drives the energy down by a product of $\bigl(1-\tfrac{1}{k}\bigr)$ terms, which is bounded by an exponential $e^{-M/N}$. Hence the projection error $\epsilon_{g}$ decays exponentially in $M$.
@@
[…]
@@
Here, $A_{n},B_{n}$ are determined by the number of iterations $n$, the same below. We now absorb both terms into a single, nn–dependent penalty. Define $\sigma_{\min}(n)\;=\;\min_{x\in\mathcal{X}}\,{\sigma}_{n}(x),\qquad\delta_{n}\;=\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)},\qquad\tilde{\beta}_{n}\;=\;\beta_{n}+\delta_{n}.$ Since for all $x$, ${\sigma}_{n}(x)\geq\sigma_{\min}(n)>0$, it follows that $A_{n}+\beta_{n}\,B_{n}=\;\delta_{n}\,\sigma_{\min}(n)\;\leq\;\delta_{n}\,{\sigma}_{n}(x),$ and hence
@@
where $\tilde{\beta}_{n}=\;\beta_{n}+\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)}.$ This completes the single‐penalty construction.